Edward has saved $225,000 for retirement, and his account earns 6% interest compounded monthly. How much can he take out of his account each month, if he wants to be able to take withdrawals for 15 years?
Answer #1

Total Savings = $ 225000, Interest Rate = 6 % per annum or (6/12) = 0.5 % per month, Tenure = 15 years or 180 months

Let the permissible monthly withdrawals be $ M

Therefore, 225000 = M x (1/0.005) x [1-{1/(1.005)^(180)}]

M = 225000 / 118.5035 = $ 1898.6778 ~ $ 1898.68